1. The reaction of 7.20 g of carbon with excess O2 yields 9.00 g of CO2. What is the percent yield of this reaction?
2. What is the molecular formula for the hydrocarbon in the mass spectrum above?
3. If we have 20 moles of the hydrocarbon (from the mass spectrum) above, how many moles of water will be produced assuming the reaction goes to completion? The reaction occurs only in the presence of the hydrocarbon and excess oxygen.
1
Expert's answer
2019-11-12T12:01:54-0500

1. C + O2 = CO2

n(C) = m/M = 7.2g/12g/mol = 0.6 mol

m(CO2) = n*M = 0.6mol*44g/mol = 26.4 g

Percent yield = 9*100%/26.4 = 34.09%

2. You didn't provide the mass spectrum for hydrocarbon.